You have $100,000 to invest in either Stock D, Stock F, or a risk-free asset. You must invest all of your money. Your goal is to create a portfolio that has an expected return of 11.5 percent. Assume D has an expected return of 15 percent, F has an expected return of 10.9 percent, and the risk-free rate is 6 percent.

If you invest $50,000 in Stock D, how much will you invest in Stock F? (Do not round intermediate calculations and round your answer to 2 decimal places, e.g., 32.16.)

Amount of Stock F to buy            $ ________
Answer #1

Ans $ 20408.16

Let Investment in Risk Free Asset be x
Investment in F = (100000 - 50000 - x) = 50000- x
RETURN OF D + RETURN OF F + RETURN ON RISK FREE ASSET = EXPECTED RETURN
50000 * 15% + (50000 - x) * 10.9% + x * 6% = 100000 * 11.5%
7500 + 5450 - 0.109x + 0.06x = 11500
12950 -11500 = 0.049x
x = 1450 / 0.049
x = 29591.83673
Investment in F = 50000- x
=50000 - 29591.84
=20408.16